OpenStudy (anonymous):

How do I determine the acceleration at 1.0s, 3.0s, and 5.5s?

OpenStudy (anonymous):

Velocity at 1.0s is 7.5 m/s [W] at 3.0s is 15.0 m/s [W] at 5.5s is 7.5 m/s [W]

OpenStudy (anonymous):

I tried to do slope a = Vf-Vi ----- delta time = 0-7.5cm/s ----- 1s = -7.5 cm but the answer is 7.5 m/s^2 [W] D:

OpenStudy (anonymous):

acceleration =\[\Delta v / \Delta t\] (7.5 m/s - 0) / (1s - 0) = 7.5 m/s^2

OpenStudy (anonymous):

do I make my acceleration = 7.5 m/s [W] - 0 / 1s = 7.5 m/s^2 [W]?

OpenStudy (anonymous):

because then I find the correct answer.. o.o

OpenStudy (stormfire1):

(7.5 m/s - 0) / (1s - 0) = 7.5 m/s^2 (15 m/s - 7.5) / (3s - 1) = ? m/s^2 (7.5 m/s - 15) / (5.5s - 3) = ? m/s^2

OpenStudy (stormfire1):

You need to make sure you get your initial and final velocities correctly placed in the formula...once you do that these are pretty easy

OpenStudy (anonymous):

ohh, yup I made a mistake when I put my initial and final velocities >_<"

OpenStudy (stormfire1):

:)

OpenStudy (anonymous):

but oh, for the velocity at 5.5s is 7.5 m/s [W], it asks for its acceleration: a = Vf-Vi ----- delta t = 0 - (-7.5 m/s [W]) ---------------- 0.5s = 7.5 m/s [W] ----------- 0.5s = 15 m/s^2 [E]

OpenStudy (anonymous):

and 15 m/s^2 [E] is the correct answer, but I'm wondering why it is east and not west..

OpenStudy (stormfire1):

Think about it: you were going 15/ms W and went down to 7.5 m/s...you had to have a negative acceleration...which would be in the opposite direction

OpenStudy (anonymous):

ohhhh!!!